1) A person deposits 1000 USD into an account that yields 9 percent interest compounded annually. Let An denote the amount of money in the account after n years.
(a) Set up a recurrence relation for An.
(b) Find an explicit formula for An.
(c) How much money will be in the account after 100 years?
1
Expert's answer
2020-10-13T14:45:53-0400

(a) "A_n = 1.09 \cdot A_{n-1}"

(b) "A_n = 1000 \cdot 1.09^{n}"

(c) "A_{100} = 1000 \cdot 1.09^{100} =" "5529040.79183"